meganmoney543:

Given: 3x + y = 1. solve for y.

Convert:

3x + y = 1 -3x -3x Answer: y=-3x+1

meganmoney543:

thxx

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @meganmoney543 thxx \(\color{#0cbb34}{\text{End of Quote}}\) No problem! :)